OpenStudy (anonymous):

What is the answer to 0=t^3 - 5t, I need to find the t's.

OpenStudy (anonymous):

\[t = \sqrt{5}\]

OpenStudy (anonymous):

t( t^2 - 5) = 0 t^2 - 5 = 0 t^2 = 5 Therefore, \[t = \sqrt{5}\]

OpenStudy (anonymous):

It is asking for the intersection of x(t) = t^2 and y(t) = t^3 - 5t I got sqrt5 for the x function, but I think I need at least two t's, right?

OpenStudy (anonymous):

t( t^2 - 5) = 0 so t=0 and t^2-5=0 t^2=5 \[t=\pm \sqrt{5}\]

OpenStudy (anonymous):

ok, thank you